Demand for Orange Juice is given as
Qd = 5000 – 2500 P + 1200 I + 650 E – 255 Ps
Suppose Income is I = Rs.500, Expectations E = 55, and Price of Ps = Rs 25.
a. Find the Demand Equation.
b. Using the demand function from part a.,
Calculate Elasticity of Demand for price range of Rs.125 and Rs.155.
c. What will be the ‘Price Elasticity of Demand’ at P = Rs.125?
d. Interpret the Elasticity of Demand calculated in (C) above.
1
Expert's answer
2020-11-26T11:47:31-0500

a) "Qd=5000-2500P+1200*500+650*55-255*25=634,375-2500P"

b) "Qd(P=125)=634,375-2500*125=321,875"

"Qd(P=155)=634,375-2500*155=246,875"

"El=\frac{321,875-246,875}{125-155}=-2,500"

c) "El=\frac{dQ}{dP}*\frac{P}{Q}=-2,500*\frac{125}{321,875}=-0.97"

d) It means that the changes in the price on 1% lead to the changes in Q on 0.97% in the opposite direction. Price grows Q decreases and vice versa